# A Survey on LLM-as-a-Judge

### v2

@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown et al., 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024c) and GPTScore(Fu et al., 2024).
@@
In addition to providing hight-quality examples for LLMs to inference, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly including the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024d) use Chain-of-Thought(CoT)(Wei et al., 2022) to provide guidance for LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. proposes Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition and thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024b) analyzed and validated the impact of position bias on LLM-as-a-judge, and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024d) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further codifies this form-filling paradigm, efficiently output evaluation results in JSON dictionary format and obtain multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
Apart from challenges in robustness, directly outputting evaluation results by LLMs also suffer from the lack of interpretability. The meaning of evaluation results from LLM evaluators is difficult to align consistently with instructions and metrics provided in prompts. To address the challenges, CLAIR(Chan et al., 2023) requires LLMs to output evaluation scores between 0-100 simultaneously with relevant reasons as explanations in JSON format, which enhancing the rationality and interpretability of the scores. FLEUR(Lee et al., 2024b) utilizes LLaVA to first provide quality scores for image captions and subsequently asks with ”Why? Tell me the reason.” for explanations with the images, captions and scores as inputs, offering a stepwise approach to provide interpretable scores.
@@ 3.2. Improvement Strategy of LLMs’ Abilities
The evaluation capabilities of LLMs is a reflection of their powerful general language understanding and generation abilities triggered by specific prompts. Methods for optimizing evaluation through prompt design, which focuses on LLMs’ in-contextual learning capabilities, require LLMs to fully comprehend the meaning of prompts and consistently follow the relevant evaluation instructions. However, even state-of-the-art LLMs like GPT4 face issues such as conceptual confusion(Hu et al., 2024), and smaller open-source LLMs which are easier to deploy as evaluators have even more limitations in their evaluation capabilities. Therefore, how to improve the evaluation capabilities of LLMs, including how to fine-tune LLMs through meta evaluation datasets and how to iteratively optimizing models based on feedback of evaluation results, is significant for improving the fundamental evaluation performance of LLM-as-a-judge.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024c) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
To better align with the requirements of evaluation tasks, many research works further transform inputs and instructions sampled from public datasets to construct more targeted training data. OffsetBias(Park et al., 2024) aims to reduce biases of LLMs by using GPT4 to generate off-topic versions of the original inputs and then having GPT-3.5 respond to the new inputs to produce bad responses. By pairing good and bad responses as training data to fine-tune the LLMs as evaluators, the biases in LLMs are significantly reduced, including length bias, concreteness bias, knowledge bias and so on. JudgeLM(Zhu et al., 2023a) enhances LLMs’ evaluation capabilities by creating different types of training data through paradigms like reference support and reference drop. CritiqueLLM(Ke et al., 2024) proposes a multi-path prompting approach, combining pointwise-to-pairwise and referenced-to-reference-free prompting strategies to restructure referenced pointwise grading data into four types, which helps create Eval-Instruct to fine-tune LLMs, addressing shortcomings in pointwise grading and pairwise comparison.
In summary, constructing meta evaluation training data targeted at specific evaluation tasks and fine-tuning LLMs can directly adjust the model’s internal parameterized knowledge and language abilities. This is the most straightforward method to improve the evaluation performance of LLM evaluators and address potential biases.
@@ 3.2.2. Iterative Optimization Based on Feedback of Evaluation Results
Fine-tuning LLMs on meta evaluation datasets give them the ability to produce evaluations which are more aligned with human preferences. However, LLM-as-a-judge may still introduce biases during evaluation process in practice, which can impact the overall evaluation quality. A natural improvement strategy is to iteratively optimize the model based on feedback of evaluation results, which mainly comes from stronger models or directly from human evaluators’ correction of the evaluation results.
A typical example is INSTRUCTSCORE(Xu et al., 2023). To improve model performance and further benefit the final quality score calculation, this score framework collects failure modes of metric outputs, query GPT-4 on each failure mode to gather automatic feedback, and finally selects explanations most aligned with human preferences to iteratively fine-tune the LLaMA model. Unlike INSTRUCTSCORE which directly optimizes the model, the LLM evaluator in JADE(Zhang et al., 2023c) relies on human judges to correct LLMs’ evaluation results and updates the most frequently corrected samples into the example sets for few-shot prompting. JADE utilizes this relatively low-cost method to achieve iterative updates of the evaluation capabilities.
Since the feedback is more closely aligned with human preferences, LLM evaluators can dynamically align with human when optimizing evaluation capabilities based on this feedback, leading to better evaluation results. This feedback-based iterative optimization strategy address the problem of models’ imperfect generalization and improve the evaluation capabilities through dynamic updates.
@@ 3.3. Optimization Strategy of Final Results
Through the optimization based on in-context learning and the model’ own capabilities, LLMs have become fairly reliable evaluators which are capable of understanding evaluation task requirements and providing rational evaluation results. However, the inherent generation randomness within the black box of LLMs still introduces significant instability to the entire evaluation pipeline, affecting the overall evaluation quality. Therefore, optimization strategies during the post-processing stage from LLM evaluators’ outputs to final evaluation results are necessary. In this survey, these optimization strategies are categorized into three types: integration of multiple evaluation results, direct optimization of LLMs’ outputs, and conversion of evaluation tasks from pointwise evaluation to pairwise comparison.
@@ 3.3.1. Integration of Multiple Evaluation Results
Integrating multiple evaluation results for the same content to obtain the final result is a common strategy in various experiments and engineering pipelines, which can reduce the impacts of accidental factors and random errors. The most basic optimization strategy is to perform multiple runs of evaluation on the same content with different hyper-parameters and settings, and then summarize these results. For example the work of Sottana et al.(Sottana et al., 2023) reduces randomness in evaluations by averaging multiple scores of the same sample. Similarly, PsychoBench(Huang et al., 2023) takes the mean and standard deviation from ten independent runs. Auto-J(Li et al., 2023b) further amplifies the differences between evaluation rounds, which combine critiques with and without scenario criteria to obtain the final results.
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and the integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023e) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022) and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2023b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
In summary, forming the final evaluation results by combining multiple rounds of evaluations or multiple LLM evaluators can reduce the random effects caused by accidental factors in a single round and reduce the potential biases of single LLM evaluator. This strategy significantly enhances the stability and reliability of the evaluation results.
@@ 3.3.2. Direct Optimization of LLMs’ Outputs
However, methods like score smoothing, which combine implicit logits and explicit outputs require the LLMs to be open-source, or to provide interfaces that allow access to token probabilities, which brings some limitations. Inspired by the work of Weng et al.(Weng et al., 2023) and Madaan et al.(Madaan et al., 2023), self-verification can be used to filter out the evaluation results without sufficient robustness. For example, TrueTeacher(Gekhman et al., 2023) applies self-verification in its evaluation of distilled data by asking the LLM evaluator for its certainty about the evaluation results after providing them, and retaining only those results that pass self-verification. Self-verification is suitable for all LLMs and require no complex computing and processing.
@@ 4.1. Basic Metric
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023d) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. In (Zeng et al., 2023), a meta-evaluation benchmark was created to evaluate the effectiveness of LLM evaluators in recognizing instruction-following outputs, focusing on only a part of the LLM-as-a-Judge pipeline.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0

### v3

@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown et al., 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes, and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024c), and GPTScore(Fu et al., 2024).
@@
In addition to providing high-quality examples for LLMs, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly include the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024d) use Chain-of-Thought(CoT)(Wei et al., 2022) to guide LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. propose Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness, and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements are specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024b) analyzed and validated the impact of position bias on LLM-as-a-judge and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024d) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further modifies this form-filling paradigm, efficiently outputs evaluation results in JSON format, and obtains multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
Apart from challenges in robustness, directly outputting evaluation results by LLMs also suffer from the lack of interpretability. The meaning of evaluation results from LLM evaluators is difficult to align consistently with instructions and metrics provided in prompts. To address the challenges, CLAIR(Chan et al., 2023) requires LLMs to output evaluation scores between 0-100 simultaneously with relevant reasons as explanations in JSON format, which enhances the rationality and interpretability of the scores. FLEUR(Lee et al., 2024) utilizes LLaVA to first provide quality scores for image captions and subsequently asks with ”Why? Tell me the reason.” for explanations with the images, captions, and scores as inputs, offering a stepwise approach to provide interpretable scores.
@@ 3.2. Improvement Strategy of LLMs’ Abilities
The evaluation capabilities of LLMs are a reflection of their powerful general language understanding and generation abilities triggered by specific prompts. Methods for optimizing evaluation through prompt design, which focuses on LLMs’ in-contextual learning capabilities, require LLMs to fully comprehend the meaning of prompts and consistently follow the relevant evaluation instructions. However, even state-of-the-art LLMs like GPT-4 face issues such as conceptual confusion(Hu et al., 2024), and smaller open-source LLMs have even more limitations in their evaluation capabilities. Therefore, how to improve the evaluation capabilities of LLMs, including how to fine-tune LLMs through meta-evaluation datasets and how to iteratively optimize models based on the feedback of evaluation results, is significant for improving the fundamental evaluation performance of LLM-as-a-judge.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta-evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024c) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
To better align with the requirements of evaluation tasks, many research works further transform inputs and instructions sampled from public datasets to construct more targeted training data. OffsetBias(Park et al., 2024) aims to reduce biases of LLMs by using GPT4 to generate off-topic versions of the original inputs and then having GPT-3.5 respond to the new inputs to produce bad responses. By pairing good and bad responses as training data to fine-tune the LLMs as evaluators, the biases in LLMs are significantly reduced, including length bias, concreteness bias, knowledge bias, and so on. JudgeLM(Zhu et al., 2023a) enhances LLMs’ evaluation capabilities by creating different types of training data through paradigms like reference support and reference drop. CritiqueLLM(Ke et al., 2024) proposes a multi-path prompting approach, combining pointwise-to-pairwise and referenced-to-reference-free prompting strategies to restructure referenced pointwise grading data into four types, which helps create Eval-Instruct to fine-tune LLMs, addressing shortcomings in pointwise grading and pairwise comparison.
In summary, constructing meta-evaluation training data targeted at specific evaluation tasks and fine-tuning LLMs can directly adjust the model’s internal parameterized knowledge and language abilities. This is the most straightforward method to improve the evaluation performance of LLM evaluators and address potential biases.
@@ 3.2.2. Iterative Optimization Based on Feedback of Evaluation Results
Fine-tuning LLMs on meta-evaluation datasets gives them the ability to produce evaluations that are more aligned with human preferences. However, LLM-as-a-judge may still introduce biases during the evaluation process in practice, which can impact the overall evaluation quality. A natural improvement strategy is to iteratively optimize the model based on the feedback of evaluation results, which mainly comes from stronger models or directly from human evaluators’ correction of the evaluation results.
A typical example is INSTRUCTSCORE(Xu et al., 2023). To improve model performance and further benefit the final quality score calculation, this scoring framework collects failure modes of metric outputs, queries GPT-4 on each failure mode to gather automatic feedback, and finally selects explanations most aligned with human preferences to iteratively fine-tune the LLaMA model. Unlike INSTRUCTSCORE which directly optimizes the model, the LLM evaluator in JADE(Zhang et al., 2023b) relies on human judges to correct LLMs’ evaluation results and updates the most frequently corrected samples into the example sets for few-shot prompting. JADE utilizes this relatively low-cost method to achieve iterative updates of the evaluation capabilities.
Since the feedback is more closely aligned with human preferences, LLM evaluators can dynamically align with humans when optimizing evaluation capabilities based on this feedback, leading to better evaluation results. This feedback-based iterative optimization strategy addresses the problem of models’ imperfect generalization and improves the evaluation capabilities through dynamic updates.
@@ 3.3. Optimization Strategy of Final Results
Through optimization based on in-context learning and the model’ own capabilities, LLMs have become fairly reliable evaluators that are capable of understanding evaluation task requirements and providing rational evaluation results. However, the inherent generation randomness within the black box of LLMs still introduces significant instability to the entire evaluation pipeline, affecting the overall evaluation quality. Therefore, optimization strategies during the post-processing stage from LLM evaluators’ outputs to final evaluation results are necessary. In this survey, these optimization strategies are categorized into three types: integration of multiple evaluation results, direct optimization of LLMs’ outputs, and conversion of evaluation tasks from pointwise evaluation to pairwise comparison.
@@ 3.3.1. Integration of Multiple Evaluation Results
Integrating multiple evaluation results for the same content to obtain the final result is a common strategy in various experiments and engineering pipelines, which can reduce the impacts of accidental factors and random errors. The most basic optimization strategy is to perform multiple runs of evaluation on the same content with different hyper-parameters and settings, and then summarize these results. For example, the work of Sottana et al.(Sottana et al., 2023) reduces randomness in evaluations by averaging multiple scores of the same sample. Similarly, PsychoBench(Huang et al., 2023) takes the mean and standard deviation from ten independent runs. Auto-J(Li et al., 2023b) further amplifies the differences between evaluation rounds, which combine critiques with and without scenario criteria to obtain the final results.
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023d) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022), and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2023b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
In summary, forming the final evaluation results by combining multiple rounds of evaluations or multiple LLM evaluators can reduce the random effects caused by accidental factors in a single round and reduce the potential biases of a single LLM evaluator. This strategy significantly enhances the stability and reliability of the evaluation results.
@@ 3.3.2. Direct Optimization of LLMs’ Outputs
However, methods like score smoothing, which combine implicit logits and explicit outputs require the LLMs to be open-source or to provide interfaces that allow access to token probabilities, which brings some limitations. Inspired by the work of Weng et al.(Weng et al., 2023) and Madaan et al.(Madaan et al., 2023), self-verification can be used to filter out the evaluation results without sufficient robustness. For example, TrueTeacher(Gekhman et al., 2023) applies self-verification in its evaluation of distilled data by asking the LLM evaluator for its certainty about the evaluation results after providing them and retaining only those results that pass self-verification. Self-verification is suitable for all LLMs and requires no complex computing and processing.
@@ 4.1. Basic Metric
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023c) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. In (Zeng et al., 2023), a meta-evaluation benchmark was created to evaluate the effectiveness of LLM evaluators in recognizing instruction-following outputs, focusing on only a part of the LLM-as-a-Judge pipeline.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023c) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0